OpenStudy (anonymous):

Solve by elimination method x+4y=19 -x+7=14

OpenStudy (lalaly):

thers a y missing?

OpenStudy (lalaly):

in the second equation

OpenStudy (asadkarim7):

elimanate x 7 may be 7y so 11y = 19 + 14 11y = 33 y = 33/ 11 y = 3 substitute x+ 4 ( 3) = 19 x + 12 = 19 x = 19 - 12 x = 7

OpenStudy (anonymous):

I apologize my computer is freaking out. Its x+4y=19 and -x+7y=14

OpenStudy (anonymous):

there should be two numbers in the answer though right? (x,y) but with numbers in place of the variables is that correct?